OpenStudy (anonymous):

Find the centroid of the region bounded by the given curves. y = 6 sin 4x, y = 6 cos 4x, x = 0, x = pi/16

OpenStudy (tkhunny):

What have you tried? Did you set up the Three Integrals?

OpenStudy (anonymous):

yeah i did. I worked it out and entered my answer but it was wrong.

OpenStudy (tkhunny):

Okay, what were your integrals?

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/16} 6 \cos 4x - 6 \sin 4x dx \]

OpenStudy (tkhunny):

Okay, there's the base. I get numerically 0.621 or so. How about the other two?

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/16} 1/2* (6\cos 4x)^2 - (6\sin 4x)^2\]

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/16} x* 6 \cos 4x - 6 \sin 4x\]

OpenStudy (tkhunny):

Hmmm... That's not quite it. The first - Maybe you just didn't put the giant parentheses around the two squared monstrosities? The second - I hope you mean \(x\cdot\left(6\cos(4x) - 6\sin(4x)\right)\). Parentheses make all the difference.

OpenStudy (anonymous):

yeah thats what I meant

OpenStudy (tkhunny):

How about the first one - the one with the squares? Both with 1/2?

OpenStudy (anonymous):

\[Its \int\limits_{0}^{\pi/16} 1/A* 1/2* (6 \cos(4x)-6 \sin(4x))^2 dx \]

OpenStudy (tkhunny):

That's no good. Should be \(½6\cos^{2}(4x) - ½6\sin^{2}(4x)\)

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

so for my x bar- I have \[((3/32) \sqrt{2}\pi-4)/(3/2)(\sqrt{2}-1)\]

OpenStudy (anonymous):

And my y bar is \[(3/8) /(3/2) (\sqrt{2}-1)\]

OpenStudy (tkhunny):

Well, that's ugly enough. xbar is no good. Why is it negative? ybar also no good. Something going wrong in there.

OpenStudy (anonymous):

when I get my Mx and My. what do i put it over?

OpenStudy (tkhunny):

The denominator is just the area of the region. Should be: \(\dfrac{3}{32}\left(\pi\sqrt{2} - 4\right)\) Looks like your close. Better algebra may be needed.

OpenStudy (tkhunny):

For the other, you should get simply 9/4.

OpenStudy (tkhunny):

The region area is just \(\dfrac{3}{2}\left(\sqrt{2} - 1\right)\)

OpenStudy (tkhunny):

Final Answer: \(\left(\dfrac{\pi\sqrt{2}-4}{16\cdot(\sqrt{2} - 1)},\dfrac{3}{2\cdot(\sqrt{2} - 1)}\right)\)